OpenStudy (anonymous):

mc115-1.jpg When NB = 12 ft, find the area of the region that is NOT shaded. Round to the nearest tenth. 150.7 ft2 88.3 ft2 452.2 ft2 363.9 ft2

OpenStudy (anonymous):

can you post the image that goes with this?

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

Hope it helps.

OpenStudy (anonymous):

the answer given by YZY is correct in short form full circle area =pi*r^2=452 minus sector which is a fraction of a full circle = 120/360*450=150 add on the triangle area 1/2 base *height=62.3 452-150.8+62.3=363.5